OpenStudy (shamil98):

integrate ∫(4-x)^11 dx u = 4-x ∫(4-x)^11 dx = ∫u^11 du = u^12 / 12 + C rite?

OpenStudy (luigi0210):

Yes, now sub back in

OpenStudy (luigi0210):

I think.. >.>

OpenStudy (shamil98):

= (4-x)^12 / 12 + C

OpenStudy (shamil98):

i looked at the solution it says that it is: - ∫u^11 du why is that?

OpenStudy (bahrom7893):

what's du?

OpenStudy (bahrom7893):

You're wrong shamil.. u=4-x du=-dx

OpenStudy (bahrom7893):

The integral becomes: -Int(u^11du)

OpenStudy (shamil98):

me no comprende, why though? o.o

OpenStudy (bahrom7893):

u = 4-x right?

OpenStudy (shamil98):

yeh

OpenStudy (bahrom7893):

u = 4x^0 - x^1 du/dx = 0*4x^(-1) - 1*x^0

OpenStudy (bahrom7893):

du/dx = -1

OpenStudy (luigi0210):

Screw it, bah's got it.

OpenStudy (shamil98):

oh

OpenStudy (bahrom7893):

du=-dx

OpenStudy (shamil98):

ty bah

OpenStudy (shamil98):

= -u^12 / 12 + C = -(4-x)^12/12 + C